# Federated Joint Learning for Domain and Class Generalization

### v1

@@ 1 Introduction
Recent advancements in visual-language models [9, 11], epitomized by CLIP [14], have demonstrated robust feature extraction through extensive pretraining on large-scale image-text pairs, achieving competitive performance across a multitude of downstream tasks [22, 24, 16, 18] such as image classification and text-to-image retrieval. Despite the impressive capabilities of CLIP, fine-tuning this model to improve performance on downstream tasks is frequently constrained by the substantial computational demands due to its large number of parameters. As a result, recent research has increasingly focused on developing efficient fine-tuning methodologies [8] to address this challenge.
@@
Inspired by previous work [2, 12, 19, 17],we propose FedDCG, a novel approach that addresses both class and domain generalization in federated learning setting based on decoupling concept. Specifically, during the training period, we adopt Class-Specific Domain-Grouping Collaborative Training Strategy to train the model to solve class generalization and domain generalization issues. In simple terms, we group domains and train a class-generalized network within each group to avoid decision boundary confusion in classes and domains. During inference, we aggregate the class-generalized results based on domain similarity, thereby integrating knowledge from both class and domain generalization. Our contributions can be summarized as follows:

### v2

@@ 1 Introduction
Recent advancements in visual-language models [9, 11, 13, 23], epitomized by CLIP [16], have demonstrated robust feature extraction through extensive pretraining on large-scale image-text pairs, achieving competitive performance across a multitude of downstream tasks [28, 30, 19, 24] such as image classification and text-to-image retrieval. Despite the impressive capabilities of CLIP, fine-tuning this model to improve performance on downstream tasks is frequently constrained by the substantial computational demands due to its large number of parameters. As a result, recent research has increasingly focused on developing efficient fine-tuning methodologies [8] to address this challenge.
@@
Inspired by previous work [2, 14, 25, 20, 18, 21, 12, 22],we propose FedDCG, a novel approach that addresses both class and domain generalization in federated learning setting based on decoupling concept. Specifically, during the training period, we adopt Class-Specific Domain-Grouping Collaborative Training Strategy to train the model to solve class generalization and domain generalization issues. In simple terms, we group domains and train a class-generalized network within each group to avoid decision boundary confusion in classes and domains. During inference, we aggregate the class-generalized results based on domain similarity, thereby integrating knowledge from both class and domain generalization. Our contributions can be summarized as follows: